2A firm manufactures two products A and B on which the profits earned per unit are Rs. 3 and Rs.4 respectively. Each product is processed on two machines M and M Product Arequires one minute of processing time on M, and two minutes on M, while requires one minute on M, and one minute on M, Machine M is available for not more than 7 hours 30 minutes while machine M is available for 10 hours during any working day. Find the number of units of products A and B to be manufactured to get maximum profit. Formulate the above as a LPP and solve by graphical method.


1
Expert's answer
2020-12-17T07:10:11-0500

Let "x" units of firm A and "y" units of firm B were manufactured: "x\geq0, y\geq0."

Given information can be tabulated as


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Time\ of\ M_1 & Time\ of\ M_2 \\ \hline
 Product\ A\ (x) & 1\min & 2\min \\
 Product\ B\ (y) & 1\min & 1\min \\
 \hdashline
 Availability & 450\min & 600\min
\end{array}"

The constrants are


"x+y\leq450, 2x+y\leq600"

Total profit "Z=3x+4y" which is to be maximized.

The objective function: "Max\ Z=3x+4y"

Subjective Constraints

"x+y\leq450"

"x+2y\leq600"

"x\geq0, y\geq0"

Horizontal "(x)"axis: Product A , Vertical "(y)" axis: Product B

Constraint No. 1: "x+y\leq450" Converting into equality: 

"x+y=450"

The two points which make the constraint line are: "(0,450)" and "(450, 0)"


Constraint No. 2: "x+2y\leq600" Converting into equality: 

"x+2y=600"

The two points which make the constraint line are: "(0,300)" and "(600, 0)"

Each constraint will be represented by a single straight line on the graph. There are two constraints, hence there will be two straight lines.

The feasible region determined by the system of constraints


"\begin{matrix}
 x+y\leq450 \\
 x+2y\leq600\\
 x\geq 0 \\
y\geq0
\end{matrix}"

is as follows



The corner points are "O(0,0), D(0,300),E(300,150),A(450,0)."

The values of "Z" at these corner points are as follows


"\begin{matrix}
 Corner\ point & Z=3x+4y \\
 O & 0 \\
 D & 1200\\
 E & 1500\\
 A & 1350
\end{matrix}"

The maximum value of "Z" is "Rs1500" which is attained at "E(300,150)."


The maximum profit is "Rs1500" obtained when 300 units of firm A and 150 products of firm B are manufactured.